OpenStudy (goformit100):

A stone is dropped into a quiet lake and waves move in circles at a speed of 4cm per second. At the instant, when the radius of the circular wave is 10 cm, how fast is the enclosed area increasing?

OpenStudy (goformit100):

@Noemi95 @Shannon20150 @9 @terenzreignz

terenzreignz (terenzreignz):

IF the waves move at 4cm per second, then the radius of the circles also increase 4cm per second... namely \[\Large \frac{dr}{dt}=4\]

OpenStudy (goformit100):

Thank you sir

terenzreignz (terenzreignz):

Enclosed area = \[\huge A=\pi r^2\] We want... \[\huge \frac{dA}{dt}=\frac{dA}{dr}\times \frac{dr}{dt}\]

OpenStudy (agent0smith):

You can also just differentiate A=pi r^2 implicitly w.r.t. t since you know dr/dt. \[\Large A=\pi r^2\] \[\Large \frac{ dA }{ dt} = 2 \pi r \frac{ dr }{ dt}\]